OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)? (1 point)

OpenStudy (anonymous):

A.) 1 B.) 2 C.) -1 D.)-2

OpenStudy (anonymous):

\[\frac{-4-1}{3-(-2)}\] is a start

OpenStudy (anonymous):

compare how much does x change and how much does y chang at the same time the equation above does just that: \[\frac{ y1 - y2 }{ x1 - x2 }\]

OpenStudy (anonymous):

Go from this way : y1-y2/x1-x2 = m The m is that the slope ! Now get this formula : y-y1=m(x-x1)

OpenStudy (anonymous):

In this question we have this : -4-1/3+2=-5/5=-1=m y+4=-1(x+3) y+4=-x-3 y=-x-7 WE always should have : y=mx+d

OpenStudy (anonymous):

so would it be -2?

OpenStudy (anonymous):

(3, -4) and (-2, 1): x1=3, y1=-4 and x2=-2, y2=1 y1--y2 = (-4) -- 1 = -5 x1--x2 = 3 -- (-2) = 5 -5/5 = -1

OpenStudy (anonymous):

does it make sense? well if we start at x=-2, we have y=1. (that is one of the points (-2, 1) ok, now we go five to the right - we end up at the second point (3, -4) because x=-2 +5 travelled lands us at x=3 and what happend to y in the meantime? it went from (-2, 1) to (3, -4). change in y was -5, when we did go 5 x to the right. Yes, the slope is -1.

OpenStudy (anonymous):

when I compare Y in the second last paragraph I always look at the second component so I compare (x, 1) and (x, -4)